Let p= sin t i + cost j + tk, What is |dp/dt|?
a.\\(\\sqrt(7)\\)
b.\\(\\sqrt(2)\\)
c.\\(\\sqrt(3)\\)
d.\\(\\sqrt(5)\\)
1
Expert's answer
2020-03-31T15:56:35-0400

"\frac{dp}{dt}=costi-sintj+k"

"|\frac{dp}{dt}|=\sqrt{cos^2t+sin^2t+1}=\sqrt{2}"

Var. b.